# IterResearch: Rethinking Long-Horizon Agents via Markovian State Reconstruction

### v1

@@ Abstract
Recent advances in deep-research agents have shown promise for autonomous knowledge construction through dynamic reasoning over external sources. However, existing approaches rely on a mono-contextual paradigm that accumulates all information in a single, expanding context window, leading to context suffocation and noise contamination that limit their effectiveness on long-horizon tasks. We introduce IterResearch, a novel iterative deep-research paradigm that reformulates long-horizon research as a Markov Decision Process with strategic workspace reconstruction. By maintaining an evolving report as memory and periodically synthesizing insights, our approach preserves consistent reasoning capacity across arbitrary exploration depths. We further develop Efficiency-Aware Policy Optimization (EAPO), a reinforcement learning framework that incentivizes efficient exploration through geometric reward discounting and enables stable distributed training via adaptive downsampling. Extensive experiments demonstrate that IterResearch achieves substantial improvements over existing open-source agents with average +14.5pp across six benchmarks and narrows the gap with frontier proprietary systems. Remarkably, our paradigm exhibits unprecedented interaction scaling, extending to 2048 interactions with dramatic performance gains (from 3.5% to 42.5%), and serves as an effective prompting strategy, improving frontier models by up to 19.2pp over ReAct on long-horizon tasks. These findings position IterResearch as a versatile solution for long-horizon reasoning, effective both as a trained agent and as a prompting paradigm for frontier models.
@@ 1 Introduction
Recent advances in deep-research agents represent a transformative shift for Large Language Models (LLMs), moving beyond passive knowledge acquisition from the model itself towards autonomous agents that construct knowledge through dynamic reasoning over external sources [dr, google_dr, grok3, Perplexity, claude_deep_research, kimi-researcher]. These frontier proprietary systems have demonstrated remarkable performance on long-horizon tasks that require sustained reasoning and information-seeking capabilities over extended interactions.
@@
When tackling long-horizon tasks, recent works [chen2025cpo, song2025r1, zheng2025deepresearcher, jin2025search, Li2025webthinker, li2025websailor, tao2025webshaper] typically append all retrieved information and intermediate reasoning steps to a single, continuously expanding context window, which we term the mono-contextual paradigm. While straightforward to implement, this paradigm fundamentally undermines the sustained reasoning capabilities required for long-horizon tasks: (1) context suffocation: as the context window fills with all prior interactions, the available space for model reasoning progressively shrinks, forcing increasingly constrained responses that ultimately degrade into premature or superficial conclusions. (2) noise contamination: irrelevant information from web searches and early exploration errors become permanently embedded in the context, creating cascading interference that dilutes signal quality throughout the entire reasoning process.
@@
To address these limitations, we introduce IterResearch, a novel Iterative Deep-Research Paradigm that fundamentally reimagines how autonomous agents maintain sustained reasoning capacity in long-horizon scenarios. Our key insight is that effective long-horizon research requires periodic synthesis and strategic forgetting—capabilities absent in current mono-contextual approaches. Specifically, we extend the Markov Decision Process (MDP) framework for deep research with a distinctive state design: rather than maintaining an ever-expanding history, each state is a strategically reconstructed workspace containing only essential elements: the question, an evolving report serving as the agent’s memory, and the immediate context needed for current reasoning. This Markovian structure, where future exploration depends only on the current reconstructed state rather than the entire history, enables the agent to maintain consistent reasoning capacity across arbitrary exploration depths while naturally circumventing the degradation that plagues mono-contextual approaches.
@@
To fully realize this paradigm’s potential, we develop Efficiency-Aware Policy Optimization (EAPO), a reinforcement learning framework specifically designed for training IterResearch. EAPO addresses two critical challenges unique to our iterative paradigm: First, recognizing that not all successful trajectories are equally valuable, we introduce efficiency-aware rewards that geometrically discount based on trajectory length—agents reaching correct conclusions through concise, focused exploration receive higher rewards than those requiring extensive iterations. Second, since our paradigm naturally decomposes trajectories into independent training samples at each round, we employ adaptive downsampling to handle the variable sample counts based on data-parallel size, ensuring stable distributed training while preserving over 99% of training data.
@@
Extensive experiments demonstrate that IterResearch significantly outperforms existing open-source agents, achieving an average improvement of 14.5 percentage points (pp) across six challenging benchmarks. More remarkably, IterResearch narrows the performance gap with frontier proprietary systems, even surpassing some on these benchmarks. Furthermore, our work reveals three fundamental insights about deep-research agents. First, our iterative paradigm unlocks extreme interaction scaling—a capability theoretically extensible to infinite depths yet structurally infeasible for current mono-contextual approaches. To our knowledge, we are the first to successfully extend agents to 2048 interactions with only 40K context length, exhibiting dramatic performance improvements (3.5% → 42.5%) as maximum interactions increase from 2 to 2048, suggesting that the perceived difficulty of long-horizon tasks may stem from insufficient exploration capacity. Second, we observe cross-paradigm knowledge transfer: trajectories generated by IterResearch significantly enhance mono-contextual agents, demonstrating that our paradigm induces superior exploration behaviors that create high-quality training signals transferable even across paradigmatically different approaches. Third, our iterative paradigm serves as an effective prompting strategy: without any training, simply applying it to frontier models yields substantial improvements over the standard mono-contextual approach, ReAct [yao2023react], particularly on long-horizon tasks (+12.7-19.2pp on BrowseComp), revealing that IterResearch offers a model-agnostic solution to long-horizon reasoning. These results confirm the effectiveness of our iterative paradigm in enabling both deeper exploration and higher-quality reasoning in long-horizon scenarios.
@@
  • We propose IterResearch, a novel iterative deep-research paradigm that reformulates long-horizon research as an MDP with strategic workspace reconstruction, maintaining sustained reasoning capacity through periodic synthesis and an evolving report memory—eliminating the context suffocation and noise contamination that plague mono-contextual approaches.
  • We develop Efficiency-Aware Policy Optimization (EAPO) with geometric discounted rewards that incentivize efficient exploration and adaptive downsampling for stable distributed training, enabling effective learning from our paradigm’s unique trajectory structure.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG is a crucial approach to overcome knowledge limitations of large language models (LLMs) by integrating external information sources [nakano2021webgpt, yu2024rankrag, AsaiWWSH24, wei2025instructrag, chen2025cpo, jin2025search, song2025r1, zheng2025deepresearcher]. However, traditional RAG methods are typically confined to static retrieval environments, such as Wikipedia, with limited exploration spaces, making them inadequate for complex, long-horizon reasoning tasks that require dynamic information gathering.
Deep Research. Recent advances in deep research [dr, google_dr, grok3, kimi-researcher] have transcended RAG’s limitations by deploying autonomous agents in real-world environments, demonstrating remarkable capabilities in navigating complex web environments and synthesizing information from diverse sources. However, existing open-source methods [li2025search, Li2025webthinker, tao2025webshaper, li2025websailor] predominantly adopt a mono-contextual paradigm, continuously appending all retrieved information and reasoning steps to a single expanding context. This linear accumulation leads to progressive workspace suffocation and irreversible noise contamination, limiting their effectiveness in long-horizon tasks. In contrast, our IterResearch reimagines deep research by formalizing it as a Markov Decision Process with workspace reconstruction mechanism, eliminating accumulation-induced degradation and enabling sustained reasoning capacity at arbitrary research depths—a critical advantage absent in existing approaches.
@@ 3 Methodology
In this section, we detail IterResearch, which extends the Markov Decision Process framework to deep research through strategic workspace reconstruction (§3.1), as illustrated in Figure 2. Then, we further introduce Efficiency-Aware Policy Optimization for training (§3.2).
Figure 2: (Top) The mono-contextual approach linearly accumulates all information into a single, ever-expanding context, leading to context suffocation and noise contamination. (Bottom) IterResearch models deep research as an extended MDP with workspace reconstruction. Each round begins with a reconstructed workspace $s_{t}$ containing the question, an evolving report $\mathcal{M}_{t}$, and immediate context. The agent generates structured decisions $d_{t}=$ (Think, Report, Action) and interacts with environment $\mathcal{E}$. The transition function $\mathcal{T}$ reconstructs the workspace, maintaining the Markov property while preventing context bloat and enabling sustained reasoning and information-seeking.
@@ 3.1 Iterative Deep-Research Paradigm
3.1.1 Markov Decision Process Formulation
We model IterResearch as an extended Markov Decision Process defined by the tuple $\langle\mathcal{S},\mathcal{D},\mathcal{E},\mathcal{T}\rangle$, where the agent conducts research through iterative rounds of exploration and synthesis to enable unbounded exploration.
  • State Space $\mathcal{S}$: Each state $s_{t}=(q,\mathcal{M}_{t},\{a_{t-1},\text{TR}_{t-1}\})$ represents the agent’s workspace, comprising question $q$, an evolving report $\mathcal{M}_{t}$ that compresses all critical findings from previous rounds, and the immediate context (action $a_{t-1}$ and tool response $\text{TR}_{t-1}$) from last interaction.
  • Decision Space $\mathcal{D}$: At each state $s_{t}$, the agent generates a structured decision $d_{t}=(\text{Think}_{t},\mathcal{M}_{t+1},a_{t})$ where: (1) Think: Reasoning about current progress and identifying information gaps. (2) Report ($\mathcal{M}_{t+1}$): Updated report serving as the agent’s compressed memory, incorporating new findings from $\text{TR}_{t-1}$ while preserving essential insights from $\mathcal{M}_{t}$ and filtering noise. (3) Action: The agent’s next operation, which can be either a tool call to gather information or a final answer when the agent determines it can adequately address the question.
  • Environment $\mathcal{E}$: External tools (Google Search, Google Scholar, Web Browser, Python) that return responses $\text{TR}_{t}=\mathcal{E}(a_{t})$ containing requested information or computation results.
  • Transition Function $\mathcal{T}$: Deterministically maps ($s_{t}\xrightarrow{d_{t},\text{TR}_{t}}s_{t+1}$) current state, decision, and tool response to the next state. Unlike mono-contextual approaches that accumulate context, we reconstruct the workspace, maintaining only the question $q$, agent-updated report $\mathcal{M}_{t+1}$, and latest interaction $\{a_{t},\text{TR}_{t}\}$, preventing context blowup.
@@ 3.1.1 Markov Decision Process Formulation
$\begin{cases}\text{Decision:}&d_{t}=\pi(s_{t})=(\text{Think}_{t},\mathcal{M}_{t+1},a_{t})\\[3.0pt] \text{Transition:}&s_{t+1}=\mathcal{T}(s_{t},d_{t},\mathcal{E}(a_{t}))=(q,\mathcal{M}_{t+1},\{a_{t},\text{TR}_{t}\})\end{cases}$ (1)
where $\text{TR}_{t}=\mathcal{E}(a_{t})$, initial state $s_{0}=(q,\mathcal{M}_{0},\emptyset)$ with empty report $\mathcal{M}_{0}$. The iterative process generates a trajectory $\tau=\{(s_{0},d_{0},\text{TR}_{0}),(s_{1},d_{1},\text{TR}_{1}),\ldots,(s_{T},d_{T})\}$ terminating when $a_{T}=\texttt{answer}$. Unlike mono-contextual approaches where context grows linearly with trajectory length, our workspace reconstruction maintains bounded memory footprint—the report $\mathcal{M}_{t}$ synthesizes findings rather than accumulating raw observations, enabling sustained reasoning quality over extended research trajectories.
3.1.2 Markovian Workspace Reconstruction
The cornerstone of our paradigm is workspace reconstruction, which fundamentally departs from traditional linear accumulation approaches [Li2025webthinker, li2025websailor, tao2025webshaper]. While existing methods suffer from $O(t)$ context growth leading to inevitable performance degradation, we introduce a principled reconstruction mechanism that maintains bounded workspace complexity while preserving complete task-relevant information through selective compression.
@@ 3.1.2 Markovian Workspace Reconstruction
$\underbrace{s_{t}^{\text{mono}}=[q,a_{0},\text{TR}_{0},...,a_{t-1},\text{TR}_{t-1}]}_{\text{Mono-contextual (ReAct): }{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{O}(t)\text{ growth}}}\quad\text{vs.}\quad\underbrace{s_{t}^{\text{iter}}=(q,\mathcal{M}_{t},\{a_{t-1},\text{TR}_{t-1}\})}_{\text{IterResearch (Ours): }{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}(1)\text{ constant}}}$ (2)
Through the markovian workspace, the agent maintains consistent reasoning capacity throughout the research process, avoiding the performance degradation that inevitably occurs when context windows approach their limits. Furthermore, through end-to-end training (§3.2), the agent progressively learns to synthesize reports that effectively filter noise and preserve essential information. Thus, irrelevant information or errors from early rounds cannot directly propagate to future decisions—they must first pass through the agent’s synthesis to be incorporated into the report. This selective retention ensures that the Markov property holds: the current state $s_{t+1}$ contains all decision-relevant information, making the full history unnecessary for optimal decision-making. The transformative impact of this design manifests in interaction scaling. While mono-contextual approaches typically fail or degrade severely beyond dozens of interactions due to context limitations, our IterResearch enables theoretically unbounded exploration, sustaining consistent reasoning quality at arbitrary depths. This scaling capability, empirically validated through experiments with up to 2048 interactions (§ 4.4), fundamentally expands the scope of problems that deep-research agents can tackle.

### v2

# IterResearch: Rethinking Long-Horizon Agents with Interaction Scaling

@@ Abstract
Recent advances in deep-research agents have shown promise for autonomous knowledge construction through dynamic reasoning over external sources. However, existing approaches rely on a mono-contextual paradigm that accumulates all information in a single, expanding context window, leading to context suffocation and noise contamination that limit their effectiveness on long-horizon tasks. We introduce IterResearch, a novel iterative deep-research paradigm that revisits long-horizon research through the lens of Interaction Scaling. Instead of relying on linear context accumulation, we adopt an MDP-inspired architecture with strategic workspace reconstruction. By maintaining an evolving report as memory and periodically synthesizing insights, our approach preserves consistent reasoning capacity across arbitrary exploration depths. To effectively train this paradigm, we employ Efficiency-Aware Policy Optimization (EAPO), a training strategy that adapts geometric reward discounting to incentivize efficient exploration and utilizes adaptive downsampling for stable distributed training. Extensive experiments demonstrate that IterResearch achieves substantial improvements over existing open-source agents with average +14.5pp across six benchmarks and narrows the gap with frontier proprietary systems. Remarkably, our paradigm exhibits unprecedented interaction scaling, extending to 2048 interactions with dramatic performance gains (from 3.5% to 42.5%), and serves as an effective prompting strategy, improving frontier models by up to 19.2pp over ReAct on long-horizon tasks. These findings position IterResearch as a versatile solution for long-horizon reasoning, effective both as a trained agent and as a prompting paradigm for frontier models.
@@ 1 Introduction
Recent advances in deep-research agents represent a transformative shift for Large Language Models (LLMs), moving beyond passive knowledge acquisition from the model itself towards autonomous agents that construct knowledge through dynamic reasoning over external sources (OpenAI, 2025a; Google, 2025; xAI, 2025; Perplexity, 2025; Anthropic, 2025; Team et al., 2025). These frontier proprietary systems have demonstrated remarkable performance on long-horizon tasks that require sustained reasoning and information-seeking capabilities over extended interactions.
@@
When tackling long-horizon tasks, recent works (Chen et al., 2025a; Song et al., 2025; Zheng et al., 2025b; Jin et al., 2025; Li et al., 2025c; a; Tao et al., 2025) typically append all retrieved information and intermediate reasoning steps to a single, continuously expanding context window, which we term the mono-contextual paradigm. While straightforward to implement, this paradigm fundamentally undermines the sustained reasoning capabilities required for long-horizon tasks: (1) context suffocation: as the context window fills with all prior interactions, the available space for model reasoning progressively shrinks, forcing increasingly constrained responses that ultimately degrade into premature or superficial conclusions. (2) noise contamination: irrelevant information from web searches and early exploration errors become permanently embedded in the context, creating cascading interference that dilutes signal quality throughout the entire reasoning process.
@@
To address these limitations, we introduce IterResearch, a novel Iterative Deep-Research Paradigm that fundamentally reimagines how autonomous agents maintain sustained reasoning capacity in long-horizon scenarios. Our key insight is that effective long-horizon research requires periodic synthesis and strategic forgetting—capabilities absent in current mono-contextual approaches. Specifically, we adopt an MDP-inspired architecture to structure deep research with a distinctive state design: rather than maintaining an ever-expanding history, each state is a strategically reconstructed workspace containing only essential elements: the question, an evolving report serving as the agent’s memory, and the immediate context needed for current reasoning. This iterative structure, where future exploration depends only on the current reconstructed state rather than the entire history, enables the agent to maintain consistent reasoning capacity across arbitrary exploration depths while naturally circumventing the degradation that plagues mono-contextual approaches.
@@
To effectively train this iterative paradigm, we employ Efficiency-Aware Policy Optimization (EAPO), a training strategy specifically designed for IterResearch. EAPO addresses two critical challenges unique to our iterative paradigm: First, recognizing that not all successful trajectories are equally valuable, we adapt standard geometric discounting to create efficiency-aware rewards that geometrically discount based on trajectory length—agents reaching correct conclusions through concise, focused exploration receive higher rewards than those requiring extensive iterations. Second, since our paradigm naturally decomposes trajectories into independent training samples at each round, we employ adaptive downsampling to handle the variable sample counts based on data-parallel size, ensuring stable distributed training while preserving over 99% of training data.
@@
Extensive experiments demonstrate that IterResearch significantly outperforms existing open-source agents, achieving an average improvement of 14.5 percentage points (pp) across six challenging benchmarks. More remarkably, IterResearch narrows the performance gap with frontier proprietary systems, even surpassing some on these benchmarks. Furthermore, our work reveals three fundamental insights about deep-research agents. First, our iterative paradigm unlocks extreme interaction scaling—a capability theoretically extensible to infinite depths yet structurally infeasible for current mono-contextual approaches. To our knowledge, we are the first to successfully extend agents to 2048 interactions with only 40K context length, exhibiting dramatic performance improvements (3.5% → 42.5%) as maximum interactions increase from 2 to 2048, suggesting that the perceived difficulty of long-horizon tasks may stem from insufficient exploration capacity. Second, we observe cross-paradigm knowledge transfer: trajectories generated by IterResearch significantly enhance mono-contextual agents, demonstrating that our paradigm induces superior exploration behaviors that create high-quality training signals transferable even across paradigmatically different approaches. Third, our iterative paradigm serves as an effective prompting strategy: without any training, simply applying it to frontier models yields substantial improvements over the standard mono-contextual approach, ReAct (Yao et al., 2023), particularly on long-horizon tasks (+12.7-19.2pp on BrowseComp), revealing that IterResearch offers a model-agnostic solution to long-horizon reasoning. These results confirm the effectiveness of our iterative paradigm in enabling both deeper exploration and higher-quality reasoning in long-horizon scenarios.
@@
  • We propose IterResearch, a novel iterative deep-research paradigm that revisits long-horizon research through an MDP-based formalism, maintaining sustained reasoning capacity through periodic synthesis and an evolving report memory—eliminating the context suffocation and noise contamination that plague mono-contextual approaches.
  • We employ Efficiency-Aware Policy Optimization (EAPO) which adapts geometric discounted rewards that incentivize efficient exploration and utilizes adaptive downsampling for stable distributed training, enabling effective learning from our paradigm’s unique trajectory structure.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG is a crucial approach to overcome knowledge limitations of large language models (LLMs) by integrating external information sources (Nakano et al., 2021; Yu et al., 2024; Asai et al., 2024; Wei et al., 2025b; Chen et al., 2025a; Jin et al., 2025; Song et al., 2025; Zheng et al., 2025b). However, traditional RAG methods are typically confined to static retrieval environments, such as Wikipedia, with limited exploration spaces, making them inadequate for complex, long-horizon reasoning tasks that require dynamic information gathering.
Deep Research. Recent advances in deep research (OpenAI, 2025a; Google, 2025; xAI, 2025; MoonshotAI, 2025) have transcended RAG’s limitations by deploying autonomous agents in real-world environments, demonstrating remarkable capabilities in navigating complex web environments and synthesizing information from diverse sources. However, existing open-source methods (Li et al., 2025b; c; Tao et al., 2025; Li et al., 2025a) predominantly adopt a mono-contextual paradigm, continuously appending all retrieved information and reasoning steps to a single expanding context. This linear accumulation leads to progressive workspace suffocation and irreversible noise contamination, limiting their effectiveness in long-horizon tasks. In contrast, our IterResearch reimagines deep research by adopting an MDP-based architecture with a iterative workspace reconstruction mechanism, eliminating accumulation-induced degradation and enabling sustained reasoning capacity at arbitrary research depths—a critical advantage absent in existing approaches.
@@ 3 Methodology
In this section, we detail IterResearch, which extends the Markov Decision Process framework to deep research through iterative workspace reconstruction (§3.1), as illustrated in Figure 2. Then, we further introduce Efficiency-Aware Policy Optimization for training (§3.2).
Figure 2: (Top) The mono-contextual approach linearly accumulates all information into a single, ever-expanding context, leading to context suffocation and noise contamination. (Bottom) IterResearch models deep research with iterative workspace reconstruction. Each round begins with a reconstructed workspace $s_{t}$ containing the question, an evolving report $\mathcal{M}_{t}$, and immediate context. The agent generates structured decisions $d_{t}=$ (Think, Report, Action) and interacts with environment $\mathcal{E}$. The transition function $\mathcal{T}$ reconstructs the workspace, maintaining an iterative state structure while preventing context bloat and enabling sustained reasoning and information-seeking.
@@ 3.1 Iterative Deep-Research Paradigm
3.1.1 MDP-inspired Formulation
To address the limitations of mono-contextual paradigms, we adopt an MDP-inspired formalism to structure long-horizon research, defined by the tuple $\langle\mathcal{S},\mathcal{D},\mathcal{E},\mathcal{T},R\rangle$. This formulation explicitly separates internal thought updates from external environmental interactions to enforce state independence.
  • State Space $\mathcal{S}$: Each state $s_{t}=(q,\mathcal{M}_{t},\{a_{t-1},\text{TR}_{t-1}\})$ represents the agent’s explicit workspace. $s_{t}$ serves as a synthesized representation of the research history, comprising: (1) the constant question $q$; (2) the evolving report $\mathcal{M}_{t}$, which acts as a compressed memory of prior validated findings; (3) the immediate context $\{a_{t-1},\text{TR}_{t-1}\}$ from the previous step.
  • Decision Space $\mathcal{D}$: At each state $s_{t}$, the policy $\pi$ generates a composite decision $d_{t}$ consisting of distinct internal thought and external actions:
    $d_{t}=[\underbrace{\text{Think}_{t},\mathcal{M}_{t+1}}_{\text{Internal Thought}},\underbrace{a_{t}}_{\text{External Action}}]\sim\pi(\cdot|s_{t})$ (1)
    Here, $\mathcal{M}_{t+1}$ represents an active memory update generated by the agent, while $a_{t}$ is the actual interaction with the environment or the final answer.
  • Environment $\mathcal{E}$ and Reward $R$: The environment executes action $a_{t}$ and returns a stochastic response $\text{TR}_{t}\sim\mathcal{E}(\cdot|a_{t})$. The reward function $R(s_{t},a_{t})$ equals $1$ if and only if the action $a_{t}$ is a terminal answer evaluated as correct by the oracle, and 0 otherwise.
  • Transition Function $\mathcal{T}$: The state transition is deterministic given the decision and environmental response. It explicitly reconstructs the workspace rather than appending to history:
    $s_{t+1}=\mathcal{T}(s_{t},d_{t},\text{TR}_{t})=(q,\mathcal{M}_{t+1},\{a_{t},\text{TR}_{t}\})$ (2)
    This reconstruction mechanism ensures that $s_{t+1}$ depends solely on the current state $s_{t}$ (via $d_{t}$) and the immediate environment feedback $\text{TR}_{t}$, strictly adhering to the MDP intuition.
@@ 3.1.1 MDP-inspired Formulation
$\begin{cases}\text{Policy Step:}&(\text{Think}_{t},\mathcal{M}_{t+1},a_{t})\sim\pi(\cdot|s_{t})\\ \text{Environment Step:}&\text{TR}_{t}\sim\mathcal{E}(\cdot|a_{t})\\ \text{State Update:}&s_{t+1}\leftarrow(q,\mathcal{M}_{t+1},\{a_{t},\text{TR}_{t}\})\end{cases}$ (3)
where $\text{TR}_{t}=\mathcal{E}(a_{t})$, initial state $s_{0}=(q,\mathcal{M}_{0},\emptyset)$ with empty report $\mathcal{M}_{0}$. The iterative process generates a trajectory $\tau=\{(s_{0},d_{0},\text{TR}_{0}),(s_{1},d_{1},\text{TR}_{1}),\ldots,(s_{T},d_{T})\}$ terminating when $a_{T}=\texttt{answer}$. Unlike mono-contextual approaches where the state dimension grows linearly ($|s_{t}|\propto t$), our formulation ensures $|s_{t}|\approx O(1)$, enabling theoretically unbounded exploration steps.
3.1.2 Iterative Workspace Reconstruction
The cornerstone of our paradigm is workspace reconstruction, which fundamentally departs from traditional linear accumulation approaches (Li et al., 2025c; a; Tao et al., 2025). While existing methods suffer from $O(t)$ context growth leading to inevitable performance degradation, we introduce a principled reconstruction mechanism that maintains bounded workspace complexity while preserving complete task-relevant information through selective compression.
@@ 3.1.2 Iterative Workspace Reconstruction
$\underbrace{s_{t}^{\text{mono}}=[q,a_{0},\text{TR}_{0},...,a_{t-1},\text{TR}_{t-1}]}_{\text{Mono-contextual: }{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{O}(t)\text{ growth}}}\quad\text{vs.}\quad\underbrace{s_{t}^{\text{iter}}=(q,\mathcal{M}_{t},\{a_{t-1},\text{TR}_{t-1}\})}_{\text{IterResearch (Ours): }{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}(1)\text{ constant}}}$ (4)
Through the iterative workspace, the agent maintains consistent reasoning capacity throughout the research process, avoiding the performance degradation that inevitably occurs when context windows approach their limits. Furthermore, through end-to-end training (§3.2), the agent progressively learns to synthesize reports that effectively filter noise and preserve essential information. Thus, irrelevant information or errors from early rounds cannot directly propagate to future decisions—they must first pass through the agent’s synthesis to be incorporated into the report. This selective retention aligns with the MDP formalism: the current state $s_{t+1}$ contains all decision-relevant information, making the full history unnecessary for optimal decision-making. The transformative impact of this design manifests in interaction scaling. While mono-contextual approaches typically fail or degrade severely beyond dozens of interactions due to context limitations, our IterResearch enables theoretically unbounded exploration, sustaining consistent reasoning quality at arbitrary depths. This scaling capability, empirically validated through experiments with up to 2048 interactions (§ 4.4), fundamentally expands the scope of problems that deep-research agents can tackle.